Let T(x,y)=6xy

be the temperature at point (x,y). A snail craws so that its position after t seconds is given by

x=1+t and y=2+13t

How fast is the temperature rising on the snail's path after 3 second?
Answer #1

Tay = 6xy x=1tt, y= 2413 € After 3 second x= 143= 4 y= 2+13 43= 41 - du dt & dy =13 dt D 24 To find 1t=3 Tony = bny Te (x,y)